OpenStudy (anonymous):

how do you simplify 2sin^2(x/2)?

OpenStudy (amistre64):

looks pretty simple as is to me :) what are ya going for?

OpenStudy (amistre64):

2(1-cos^2(x/2)) = 2 - 2cos^2(x/2) doesnt seem to help..

OpenStudy (anonymous):

i got that, but what do I do next? or is that not right

OpenStudy (amistre64):

it all depends on what your definition of "simplified" means... are we wanting only single exponents?

OpenStudy (amistre64):

err...exponents of 1?

OpenStudy (anonymous):

yes, just one term

OpenStudy (amistre64):

one term it already is, but do you mean like sin(x/2) instead of sin^2(x/2)?

OpenStudy (anonymous):

yes, i believe so

OpenStudy (amistre64):

hmmmm...... we can play around with identites some and see what happens i spose... do you recall the formula for reducing an exponent?

OpenStudy (anonymous):

nope, im so hopless haha

OpenStudy (amistre64):

cos(2a) = cos^2 - sin^2 cos(2a)= 1 - 2sin^2(2a) cos(t) -1 ------- = sin^2(t) 2 try this one

OpenStudy (amistre64):

2(cos(x/2) -1) ------------ = sin^2(x/2) 2

OpenStudy (anonymous):

okay, hold on ne second while i work it out

OpenStudy (amistre64):

.....2sin^2(x/2)

OpenStudy (anonymous):

okay, that worked. i understand that, thanks!

OpenStudy (amistre64):

that needs alittle fixing... cos(x)-1 -------- = sin^2(x/2) 2 thats what it should be before you put it in

OpenStudy (amistre64):

cos(x) - 1 = 2sin^2(x/2)

OpenStudy (anonymous):

ohh, okay, I see it now.

OpenStudy (amistre64):

better? :)

OpenStudy (anonymous):

yupp, thanks a lot!

OpenStudy (amistre64):

to double check, throw in a number and see :)